Find the x and y intercept:
x+y=6

Answer by math-vortex(648)

Hi, there--

The Problem:
Find the x and y intercept:
x+y=6

A Solution:
You have a linear equation here. By definition, the x-intercept is a coordinate pair solution 
with a y-value of 0. On a graph, this is the point where the the line crosses the x-axis. 

Similarly, the y-intercept is the coordinate pair solution with an x-value of 0. This is the point 
where the line intersects the y-axis.

Find the x-intercept:
A straightforward way to solve this problem is to substitute 0 for y in your equation.
x + y = 6
x + (0) = 6
x = 6

When y=0, x=6. The x-intecept is 6. The line x+y=6 crosses the x-axis at the point (6,0).


Find the y-intercept:
To solve this problem, substitute 0 for x in the equation.
x + y = 0
(0) + y = 6
y = 6

When x=0, y=6. The y-intercept is also 6. The line x+y=6 crosses the y-axis at the point (6,0).

NOTE: The x- and y-intercepts are not always the same number.

Hope this helps. Please email me if this is unclear or you still have questions.

Ms. Figgy
math.in.the.vortex@gmail.com